Ryan Neal bought 1,200 shares of Ford at $15.98 per share. Assume a commission of 2% of the purchase price. Ryan sells the stock for $22.25 with the same 2% commission rate. What is the gain or loss for Ryan? (Input the amount as a positive value. Round your answer to the nearest cent.)
Answer #1

Purchase cost = Price+ commission

= 15.98+2% = $16.2996

Net Proceeds from sale = 22.25 - 2% = $21.805

Gain = (Sale Price - Purchase cost)*Number of shares

= (21.805 - 16.2996)*1200

= 6,606.48